OpenStudy (anonymous):

Some numbers can be expressed as the sum of a string of consecutive positive numbers. Which numbers have this property?

OpenStudy (anonymous):

For example: \[9=2+3+4\]\[11=5+6\]\[18=3+4+5+6\]

OpenStudy (akshay_budhkar):

sum of how many consecutive numbers?

OpenStudy (anonymous):

Any amount. Try doing all the numbers from 1-20. I think you'll see something interesting.

OpenStudy (akshay_budhkar):

no no i am right..... all numbers except the power of 2!

OpenStudy (anonymous):

We have a winrar! There is some property of powers of 2 that makes them unable to be the sum of consecutive positive numbers. Bonus imaginary points if you can figure out what that property is

OpenStudy (akshay_budhkar):

lol.. i will try finding it.. wait

OpenStudy (akshay_budhkar):

give up dalvoron..:(

OpenStudy (anonymous):

Powers of two have no odd factors besides 1!

OpenStudy (akshay_budhkar):

lol we all know that! but whats the connection here?

OpenStudy (anonymous):

That's it. Any number that has an odd factor that isn't 1 can be expressed as a sum of consecutive positive numbers. Simples!